OpenStudy (anonymous):

Which isotope, X, is produced when lawrencium-256 decays by alpha emission? A: 252 over 101 Lr B: 260 over 105 Db C: 252 over 101 Md D: 256 over 101 Md

OpenStudy (anonymous):

\[\frac{ 256 }{ 103 } Lr \rightarrow X + \frac{ 4 }{ 2 } He\]

OpenStudy (anonymous):

Solve For X

OpenStudy (jfraser):

@aphex is correct, find the isotope notation that makes the equation balanced. remember to conserve total atomic number and total mass number when balancing nuclear reactions

OpenStudy (jfraser):

\[^{256}_{103}Lr \rightarrow \space ^A_ZX + \space ^4_2 He\] 256 = A + 4 and 103 = Z + 2

OpenStudy (anonymous):

So, im just guessing here, correct me if im worng. D?

OpenStudy (anonymous):

Hello?

OpenStudy (jfraser):

can 't be D, the mass number is too high

OpenStudy (jfraser):

you can't start with a mass of 256, lose 4, and still have a mass of 256

OpenStudy (anonymous):

Then its between C & A Right?

OpenStudy (jfraser):

true. What's the only difference between options A and C?

OpenStudy (anonymous):

Its C

OpenStudy (jfraser):

it is. GJ